OpenStudy (anonymous):

i dont understand this. For equivalent expressions, how do u know the expression is equivalent? If for example, 3(n-4) gets the same answer as 3n-12. does that determine its "equivalent"? or is there something more?

OpenStudy (texaschic101):

Equivalent expressions are equivalent if their values obtained by subbing the values of the variables are the same. 3(n - 4) and 3n - 12 are equivalent because the value of both expressions remains the same for any value of x.

OpenStudy (texaschic101):

oops....typo.....value of n

OpenStudy (anonymous):

ok can u give me an example of what would not be an equivalent expression

OpenStudy (texaschic101):

9n + 21 and -9n + 21....not equivalent because if you sub in 2 for n... 9(2) + 21 = 18 + 21 = 39 -9(2) + 21 = -18 + 21 = 3

OpenStudy (anonymous):

ohhh ok